OpenStudy (anonymous):

A = • A = batting average • B = times at bat multiplied by player’s seasonal average • K = hits in the new game • G = times at bat plus times at bat in a new game a. Solve A = for K. Please show your work. (6 points) b. Use your new formula and the values given in the table below to find the player’s hits in a new game (K). Round your answer to the nearest whole number. (4 points) Batter A B G K George 0.319 10.7 46 Jose 0.320 11.4 45 Dale 0.298 12.3 52 Michael 0.337 15.2 48

OpenStudy (anonymous):

I had to read this a few times to understand it, but i came up with the conclusion of 1.37 1.26 1.25 1.06 lol

OpenStudy (anonymous):

this formular is best to use A = (b+k)/g A*g = b + k K = A*g - b I hope you understand